7. Classify each variable as quantitative discrete, quantitative continuous or qualitative. A) Weight of lobsters in a restaurants tank. B) Marital status of students in our class. C) number of pairs of shoes you own 8. On a test of 80 questions, Ariel got 65 correct. What percent were correct? On a test of 20 questions. Adam got 14 correct What percent were correct? .9. On a test Roshana got 65% or 39 of problems correct. How many problems were on the test?
Answer #1

Ans:

7.

a)Quantitative Continuous

b)Qualitative

c)Quantitative Discrete

8.

i)Percentage of correct answers=(65/80)*100=81.25 %

ii)Percentage of correct answers=(14/20)*100=70 %

9)Let x=number of problems on test

Percentage of correct answer=65=(39/x)*100

0.65*x=39

x=39/0.65

x=60